Two slits separated by 1.2 mm are placed 1.40 m from a screen. Light of wavelength 480 nm and 640 nm is incident on the slits. What is the separation distance on the screen of intensity maxima of the second order?


1
Expert's answer
2021-12-20T10:30:03-0500

Find the angles for the two wavelengths:


"m\lambda_1=d\sin\theta_1,\\
m\lambda_2=d\sin\theta_2.\\\space\\
\theta_1=\arcsin\frac{m\lambda_1}{d}=0.046°,\\\space\\
\theta_2=\arcsin\frac{m\lambda_2}{d}=0.061°."

The distance between the maxima:


"x=L(\tan\theta_2-\tan\theta_1)=0.4\text{ mm}."